suppose that two objects attract each other with a gravitational force of 16 units. if the distance between the two objects is reduced by a factor of 5, then what is the new force of attraction between the two objects?


1
Expert's answer
2021-11-01T12:53:40-0400
"F_{new}=\dfrac{Gm_1m_2}{(\dfrac{d}{5})^2}=\dfrac{Gm_1m_2}{\dfrac{1}{25}d^2}=25\dfrac{Gm_1m_2}{d^2}=25F,""F_{new}=25\times16\ units=400\ units."